assuming that the heights of boys in a high school basketball tournament are normally distributed, with mean 70 inches and standard deviation 2.5 inches, how many boys in a group of 40 do we expect to be taller than 75 inches?
Answer by stanbon(75887)

Assuming that the heights of boys in a high school basketball tournament are normally distributed, with mean 70 inches and standard deviation 2.5 inches, how many boys in a group of 40 do we expect to be taller than 75 inches?
------------------
z(75) = (75-70)/2.5 = 2
---
P(a random boy is taller than 75") = P(z > 2) = 0.02275
----
Expected # of group of 40 to be over 75" = 0.02275*40 = 0.9100
Rounded up the expected number is 1.
======================================
Cheers,
Stan H.
==================